if "f(3)" =5 and "f(5)" =3, what is the form of "f(x)" by lagrange's formula


1
Expert's answer
2021-06-11T18:35:21-0400

Lagrange First Order Interpolation Formula


"f(x) = f(x_0)+(x − x_0)\dfrac{f(x_0) − f(x_1)}{x_0 − x_1}"

Given "x_0=3, f(x_0)=f(3)=5,"

"x_1=5, f(x_1)=f(5)=3"

Then


"f(x) = 5+(x − 3)\dfrac{5-3}{3-5}"

"f(x) = 5-(x \u2212 3)"

"f(x) =-x+8"